# HiGen: Hierarchy-Aware Sequence Generation for Hierarchical Text Classification

### 24 Jan 2024 v1

@@ Abstract
Hierarchical text classification (HTC) is a complex subtask under multi-label text classification, characterized by a hierarchical label taxonomy and data imbalance. The best-performing models aim to learn a static representation by combining document and hierarchical label information. However, the relevance of document sections can vary based on the hierarchy level, necessitating a dynamic document representation. To address this, we propose HiGen, a text-generation-based framework utilizing language models to encode dynamic text representations. We introduce a level-guided loss function to capture the relationship between text and label name semantics. Our approach incorporates a task-specific pretraining strategy, adapting the language model to in-domain knowledge and significantly enhancing performance for classes with limited examples. Furthermore, we present a new and valuable dataset called ENZYME, designed for HTC, which comprises articles from PubMed with the goal of predicting Enzyme Commission (EC) numbers. Through extensive experiments on the ENZYME dataset and the widely recognized WOS and NYT datasets, our methodology demonstrates superior performance, surpassing existing approaches while efficiently handling data and mitigating class imbalance. The data and code will be released publicly.
@@ 1 Introduction
Hierarchical text classification (HTC) is a task that involves categorizing text data into predefined categories organized in a hierarchical structure Banerjee et al. (2019); Wang et al. (2022a); Zhou et al. (2020). It holds great importance in various text mining applications, including scientific paper recommendation Zhang et al. (2020), semantic indexing Li et al. (2019), and online advertising Agrawal et al. (2013). HTC poses unique challenges when compared to traditional text classification, as it deals with imbalanced data distributions and complex dependencies between multiple levels of categories within the hierarchy. The hierarchical structure is represented as a directed acyclic graph (DAG), which must be encoded in the predictive model along with the text to generate the final hierarchal label. However, the data imbalance becomes more pronounced as we move down the levels, presenting a significant challenge for HTC.
@@
Inspired by the success of pretraining tasks Gururangan et al. (2020), we propose a task-specific pretraining strategy for our generative model. Our approach presents a notably simplified design in contrast to previous works which use complex encoder architectures or training paradigms [1, 2, 3, 4]. We propose a pretraining step that takes advantage of weak supervision to jointly model text representations and hierarchical label information. The encoder encodes the input document and masked hierarchical label, while the decoder regenerates the original label. The pretraining step leverages weakly labeled data from the same domain, generated using an LLM. This domain-specific knowledge, discussed in Mueller et al. (2022); Gururangan et al. (2020), proves advantageous for addressing the data imbalance challenge commonly encountered in HTC datasets.
@@
  • We present ENZYME, a dataset of 30523 full-text PubMed articles with Enzyme Commission (EC) numbers. It features a four-level single-path hierarchy, making it larger than any existing datasets for biomedical HTC.
@@ 2 Related Work
HTC is a multi-label classification problem where the classification task is performed in a hierarchical manner. Most prior work in HTC can be categorized into three broad categories: local, global and generative.
@@
Generative HTC Early works on text generation for HTC Yang et al. (2018); Risch et al. (2020) use RNN and Transformer-based seq2seq models, with dynamic document representations outperforming static encoder methods. Recent approaches Yu et al. (2022); Huang et al. (2022) propose T5-based models Raffel et al. (2020). The former addresses label inconsistency with DFS-based linearization and constrained decoding, while the latter captures dependencies using BFS-based linearization and hierarchical path-based attention.
@@ 4.2 Sequence to Sequence Modeling
Figure 2: The proposed model’s architecture consists of an encoder that takes the document and its corresponding label name as input. The decoder generates a hierarchical label output with 2 levels. To calculate the first three losses, the LM Head predicts the distribution over the vocabulary, and the hierarchy edges are considered. For the semantic loss the text and label name representations from the encoder are projected onto a shared embedding space. Positive document and label semantics are pulled together, while negatives are pushed apart. The margins $\alpha_{1}$ and $\alpha_{2}$ control the attraction between levels 1 and 2, with $\alpha_{1}>\alpha_{2}$.
@@ 4.3 Training Objective
Output Space Loss. The label hierarchy is generally represented as a DAG, where each edge signifies a parent-to-child relationship. In HTC, imposing this unidirectionality during training helps the model comprehend these hierarchical relations. To this end, we use the formulation proposed by Zhang et al. (2021) inspired by the distributional inclusion hypothesis Geffet and Dagan (2005).
@@
where $N$ is the batch size, $l$ is the number of predicted node labels and $n$ is the number of hierarchy edges. $\pi_{p}$ and $\pi_{c}$ represent the predicted probabilities for a parent label node $p$ and its child label node $c$ in the output distribution over the vocabulary. Note that penalization occurs when $\pi_{c}>\pi_{p}$ and not otherwise, preserving the unidirectionality of edges.
Token Constraint Loss. Using a model trained for open-ended text generation in HTC can lead to irrelevant node labels due to the generation of stray tokens. To address this and avoid restricting BART’s generation capabilities with a fixed vocabulary, instead, we introduce a loss function that penalizes predictions outside a designated vocabulary represented by the red dotted lines in Figure 2. This vocabulary is constructed using the label hierarchy, ensuring alignment and enhancing overall results. Putting it formally,
@@
where $\pi^{i}$ represents a vector of predicted token probabilities for a training example, $V$ is the entire vocabulary, $V^{H}$ is the desired vocabulary for $H$, and $V^{H^{\prime}}$ contains the remaining tokens. The loss function guides the model to learn the desired vocabulary’s contents, discouraging the generation of irrelevant labels and enhancing classification performance.
@@
$\mathcal{L}_{S}^{k}=max(0,\gamma_{k}^{+}-\gamma_{k}^{-}+\alpha_{k})$ (11)
@@
$min\text{ }\mathbf{\mathcal{L}_{HiGen}}=\mathcal{L}_{LM}+\lambda_{1}\mathcal{L}_{O}+\lambda_{2}\mathcal{L}_{T}+\lambda_{3}\mathcal{L}_{S}$
@@ 5.1.1 Pretraining Datasets
For the ENZYME dataset, we use the articles extracted from PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/advanced/. We randomly sample 200,000 articles along with their Enzyme Commission (EC) numbers that follow a hierarchical structure. These articles are loosely labeled as they do not reflect human-annotated EC numbers. WOS hosts a comprehensive collection of scientific articles spanning various domains of science. To produce meaningful and diverse abstracts, we employ the powerful ChatGPT model and generate $\sim$3000 abstracts for pretraining. More details are mentioned in Appendix E. As for NYT, we have access to a vast repository of articles that were not assigned to specific training, testing, or validation sets by Zhou et al. (2020). This invaluable resource enables us to utilize these uncategorized articles for our proposed pretraining task. During the process of consolidating pretraining dataset, we made sure to prevent any overlap of the pretaining data with the training, validation and testing sets.
5.1.2 ENZYME Dataset

### 22 Feb 2024 v2

@@ Abstract
Hierarchical text classification (HTC) is a complex subtask under multi-label text classification, characterized by a hierarchical label taxonomy and data imbalance. The best-performing models aim to learn a static representation by combining document and hierarchical label information. However, the relevance of document sections can vary based on the hierarchy level, necessitating a dynamic document representation. To address this, we propose HiGen, a text-generation-based framework utilizing language models to encode dynamic text representations. We introduce a level-guided loss function to capture the relationship between text and label name semantics. Our approach incorporates a task-specific pretraining strategy, adapting the language model to in-domain knowledge and significantly enhancing performance for classes with limited examples. Furthermore, we present a new and valuable dataset called ENZYME, designed for HTC, which comprises articles from PubMed with the goal of predicting Enzyme Commission (EC) numbers. Through extensive experiments on the ENZYME dataset and the widely recognized WOS and NYT datasets, our methodology demonstrates superior performance, surpassing existing approaches while efficiently handling data and mitigating class imbalance. We release our code and dataset here: https://github.com/viditjain99/HiGen.
@@ 1 Introduction
Hierarchical text classification (HTC) is a task that involves categorizing text data into predefined categories organized in a hierarchical structure Banerjee et al. (2019); Wang et al. (2022a); Zhou et al. (2020). It holds great importance in various text mining applications, including scientific paper recommendation Zhang et al. (2020); Xu et al. (2023), semantic indexing Li et al. (2019), and online advertising Agrawal et al. (2013). HTC poses unique challenges when compared to traditional text classification, as it deals with imbalanced data distributions and complex dependencies between multiple levels of categories within the hierarchy. The hierarchical structure is represented as a directed acyclic graph (DAG), which must be encoded in the predictive model along with the text to generate the final hierarchal label. However, the data imbalance becomes more pronounced as we move down the levels, presenting a significant challenge for HTC.
@@
Inspired by the success of pretraining tasks Gururangan et al. (2020), we propose a task-specific pretraining strategy for our generative model. Our approach presents a notably simplified design in contrast to previous works which use complex encoder architectures or training paradigms Zhou et al. (2020); Deng et al. (2021); Jiang et al. (2022). We propose a pretraining step that takes advantage of weak supervision to jointly model text representations and hierarchical label information. The encoder encodes the input document and masked hierarchical label, while the decoder regenerates the original label. The pretraining step leverages weakly labeled data from the same domain, generated using an LLM. This domain-specific knowledge, discussed in Mueller et al. (2022); Gururangan et al. (2020), proves advantageous for addressing the data imbalance challenge commonly encountered in HTC datasets.
@@
  • We present ENZYME, a dataset of 30,523 full-text PubMed articles with Enzyme Commission (EC) numbers. It features a four-level single-path hierarchy, making it larger than any existing datasets for biomedical HTC.
@@ 2 Related Work
HTC is a multi-label classification problem where the classification task is performed in a hierarchical manner. Most prior work in HTC can be categorized into three broad categories: local, global, and generative. In the local approach, each hierarchical level has its own classifier tailored to the unique classes at that level. Conversely, the global approach employs a unified classifier that encompasses all classes across every hierarchy level. The generative approach, a recent advancement in hierarchical text classification, capitalizes on the generative capabilities of language models to predict text labels. More details on each of the categories are explained below.
@@
Generative HTC Early works on text generation for HTC Yang et al. (2018); Risch et al. (2020) use RNN and Transformer-based seq2seq models, with dynamic document representations outperforming static encoder methods. Recent approaches Yu et al. (2022); Huang et al. (2022) propose T5-based models Raffel et al. (2020). The former addresses label inconsistency with DFS-based linearization and constrained decoding, while the latter captures dependencies using BFS-based linearization and hierarchical path-based attention. These methods overlook the dependencies between the label name from different levels and the document text. We overcome this by capturing the document-label-name dependency across different levels using the proposed level-guided semantic loss.
@@ 4.2 Sequence to Sequence Modeling
Figure 2: The proposed model’s architecture consists of an encoder that takes the document and its corresponding label name as input. The decoder generates a hierarchical label output with 2 levels. To calculate the first three losses, the LM Head predicts a distribution over the vocabulary, and the hierarchy edges are considered. For the semantic loss, the text and label name representations from the encoder are projected onto a shared embedding space. Positive document and label semantics are pulled together, while negatives are pushed apart. The margins $\alpha_{1}$ and $\alpha_{2}$ control the attraction between levels 1 and 2, with $\alpha_{1}>\alpha_{2}$.
@@ 4.3 Training Objective
Output Space Loss. The label hierarchy is generally represented as a DAG, where each edge signifies a parent-to-child relationship. In HTC, imposing this unidirectionality during training helps the model comprehend these hierarchical relations. To this end, we use the formulation proposed by Zhang et al. (2021) inspired by the distributional inclusion hypothesis (DIH) Geffet and Dagan (2005).
@@
where $N$ is the batch size, $l$ is the number of predicted node labels and $n$ is the number of hierarchy edges. $\pi_{p}$ and $\pi_{c}$ represent the predicted probabilities for a parent label node $p$ and its child label node $c$ in the output distribution over the vocabulary.
Under the DIH framework, the loss term can be precisely articulated as follows: if a document $d$ belongs to a child class $c$ with probability $\pi_{c}$, then it must belong to the parent class $p$ with a probability no less than $\pi_{c}$. For instance, if a document has a 75% chance of being labeled with "Football", the likelihood of it being assigned to the parent category "Sports" should be 75% or higher. Note that the loss term exhibits asymmetry, being non-zero when $\pi_{c}>\pi_{p}$ but zero when $\pi_{p}>\pi_{c}$. Referring to Figure 2, loss calculation involves aggregating predicted probabilities for valid tokens (indicated by red dotted lines) from the LM Head and establishing parent-child pairs using the hierarchy edges. Through the imposition of a penalty when $\pi_{c}>\pi_{p}$, the objective function guides the model to learn the correct sequence of tokens, consequently learning the unidirectionality and orientation of the hierarchy edges.
Token Constraint Loss. Using a model trained for open-ended text generation in HTC can lead to irrelevant node labels due to the generation of stray tokens (non-existent in a given hierarchy). To address this and avoid restricting BART’s generation capabilities due to a fixed vocabulary, instead, we introduce a loss function that penalizes predictions outside of a designated vocabulary represented by the red dotted lines in Figure 2. This vocabulary is constructed using the label hierarchy, ensuring alignment and enhancing overall results. Putting it formally,
@@
where $l$ is a token, $\pi^{i}$ represents a vector of predicted token probabilities for a training example by the LM Head, $V$ is the entire vocabulary, $V^{H}$ is the desired vocabulary for $H$, and $V^{H^{\prime}}$ contains the remaining tokens. The loss function guides the model to learn the desired vocabulary’s contents, discouraging the generation of irrelevant labels and enhancing classification performance.
@@
$\mathcal{L}_{S}^{k}=\max(0,\gamma_{k}^{+}-\gamma_{k}^{-}+\alpha_{k})$ (11)
@@
$\min\text{ }\mathbf{\mathcal{L}_{HiGen}}=\mathcal{L}_{LM}+\lambda_{1}\mathcal{L}_{O}+\lambda_{2}\mathcal{L}_{T}+\lambda_{3}\mathcal{L}_{S}$
@@ 5.1.1 Pretraining Datasets
For the ENZYME dataset, we use the articles extracted from PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/advanced/. We randomly sample 200,000 articles along with their Enzyme Commission (EC) numbers that follow a hierarchical structure. These articles are loosely labeled as they do not reflect human-annotated EC numbers. WOS hosts a comprehensive collection of scientific articles spanning various domains of science. To produce meaningful and diverse abstracts, we employ the powerful ChatGPT model and generate $\sim$3000 abstracts for pretraining. More details are mentioned in Appendix E. As for NYT, we have access to a vast repository of articles that were not assigned to specific training, testing, or validation sets by Zhou et al. (2020). This invaluable resource enables us to utilize these uncategorized articles for our proposed pretraining task. During the process of consolidating the pretraining dataset, we made sure to prevent any overlap of the pretraining data with the training, validation and testing sets.
@@ 6 Conclusion
Acknowledgement
This work was supported in part by NSF IIS-2008334, IIS-2106961, and CAREER IIS-2144338.